I dont understand how to solve equations like this 3/5x-7=17 when it comes to fractions I get all confused I understand problems as in 7x-5+9=21 I solve it the following 7x-5x becomes 2x so you then have 2x+9=21 then I would subtract 9 from both sides of the = sign giving me 2x+9-9=21-9 giving me 2x=12 the take 2x/2=12/2 which equals x=6 because 2*6=12 which makes x=6 but when it comes to the fraction in the problem I get all confussed as in this problem 3/5x-7=17 I dont know what to do with the fraction so if you could explain this to me I would really appreciate it very much thanks for your help we are on a block schedule so I dont have this class until Wednesday Nov 30th so I am tring o get this finished by Wednesday Thanks Again Nick
Answer by venugopalramana(3286)

SEE MY REMARKS BELOW IN CAPITAL LETTERS....
I dont understand how to solve equations like this 3/5x-7=17.... OK
when it comes to fractions I get all confused ....DONT WORRY IN A FEW MOMENTS YOUR CONFUSION WILL BE CLEARED!
I understand problems as in 7x-5X+9=21 .....GOOD
I solve it the following 7x-5x becomes 2x so you then have 2x+9=21 then I would subtract 9 from both sides of the = sign giving me 2x+9-9=21-9 giving me 2x=12 the take 2x/2=12/2 which equals x=6 because 2*6=12 which makes x=6 ......GOOD
but when it comes to the fraction in the problem I get all confussed as in this problem 3/5x-7=17....OK...YOU ARE NOT COMFORTABLE WITH FRACTIONS...SO BEST WAY IS TO GET RID OF THEM...HOW DO WE DO IT ?FIND OUT LCM OF ALL DENOMINATORS IN THE EQUATION..THEN MULTIPLY EACH AN EVERY TERM IN THE EQUATION WITH THAT...THEN THE FRACTIONS WILL DISAPPEAR!!!!
HERE WE HAVE ONLY ONE FRACTION 3/5...ITS DENOMINATOR IS 5 ..SINCE THERE ARE NO OTHER FRACTIONS LCM IS 5 ...SO WE MULTIPLY ALL TERMS WITH 5..WE GET
5*(3/5)X-5*7=5*17
3X-35=85......SEE WE GOT RID OF THE FRACTIONS.NOW THIS IS A KNOWN TERRITORY TO YOU!
3X=85+35=120
X=120/3=40

I dont know what to do with the fraction so if you could explain this to me I would really appreciate it very much thanks for your help we are on a block schedule so I dont have this class until Wednesday Nov 30th so I am tring o get this finished by Wednesday Thanks Again Nick